### v1

@@ 1. Introduction
When Question 1.1 admits an affirmative answer, we say $\mathbf{R}$ has the lifting property. There have been numerous studies on the lifting problem and its variants, for many important properties $\mathbf{R}$. For example, lifting holds for $\mathbf{R}$=“Nagata” [Marot-Nagata-Lift] and $\mathbf{R}$=“quasi-excellent” [formal-lifting-excellence-Gabber], but not for $\mathbf{R}$=“excellent” or $\mathbf{R}$=“universally catenary” [Greco-Universal-Catenary-No-Lift]. We refer the reader to [formal-lifting-excellence-Gabber, Appendix] for more information.
In previous work (see [Lyu-dual-complex-lift, §8]), the author showed that $\mathbf{R}$=“is a quotient of a Gorenstein ring” satisfies the lifting property. However, whether or not $\mathbf{R}$=“is a quotient of a Cohen–Macaulay ring” satisfies the lifting property seems to be difficult. We provide two perspectives on this question.
@@
We call a ring $R$ semi-Nagata if $R$ is Noetherian and for every finite $R$-algebra $B$ that is an integral domain, $B$ admits a finite $(S_{2})$-ification, in the sense that there is a finite inclusion of integral domains $B\to C$ so that $C$ is $(S_{2})$ and $B_{\mathfrak{p}}=C_{\mathfrak{p}}$ for all $\mathfrak{p}\in\operatorname{Spec}_{1}(B)$. We have the following main result, which the author believes to be new (Definition 7.1 and Theorems 7.8 and 7.13)
@@ 4. $(iv)$
See [EGA4_2, §7.6 and §7.7] for $(i)$$(ii)$$(iii)$, whereas $(iv)$ is trivial. In particular, from either $(iii)$ or $(iv)$ of both theorems, we have the following result, which the author also believes to be new.
@@
There are DVRs that are not Nagata, [Nagata-local, Appendix A1, Example 3]. On the other hand we have (Remark 7.2 and Corollary 7.14)
@@
To show $R^{\sigma}$ is $(S_{2})$, and to show our main Theorem 1.2, we show that for a semilocal $R$, there exists a finite subalgebra of $R^{\sigma}$ that do not have FONSIs (Theorem 6.4). After all, FONSIs are pretty rare (Remarks 4.6 and 4.7 and Lemma 6.3). The idea is inspired by a classical argument of Ratliff [Mat-CRT, §31, Lemma 4]. We use local cohomology to make a conceptual argument.
@@
Unfortunately, the author was not able to show lifting of the semi-Nagata property. However, when restricted to semilocal rings, advances in the lifting problem are made in the second part (§§812) of this article. We repeat the problem in this setting.
Question 1.6 (local lifting problem).
@@
Question 1.6 admits an affirmative answer when $\mathbf{R}$=“has $(S_{k})$ formal fibers,” where $k\geq 0$ is arbitrary, “has Cohen–Macaulay formal fibers,” “has Gorenstein formal fibers,” “has lci formal fibers,” and “is a quotient of a Cohen–Macaulay ring.”
Note that when $k=1$, this is exactly the semi-Nagata property (Theorem 1.2$(i)$).
We show the formal fiber properties in Theorem 1.7 following the argument of Nishimura [Nishimura-semilocal-lifting]. The key new input is that we find ideals that define the non-$\mathbf{P}$-locus (where $\mathbf{P}$=“$(S_{k})$,”“Cohen–Macaulay,” “Gorenstein,” or “lci”) of nice rings strictly functorially with respect to nice homomorphisms. This task was effortless for the properties considered in [Nishimura-semilocal-lifting]. We explain this in §8.
In our case, such an assignment of ideals can be found with effort. We do a basic reduction in §9, saying we just need to assign $\mathfrak{m}$-primary ideals to complete local rings $(A,\mathfrak{m})$ that are $\mathbf{P}$ exactly on the punctured spectrum, strictly functorial with respect to flat maps with $\mathbf{P}$-fibers and 00-dimensional special fiber. After the reduction we find a desired assignment for all properties but lci in §10. The case $\mathbf{R}$=“is a quotient of a Cohen–Macaulay ring” of Theorem 1.7 follows from the case $\mathbf{R}$=“has Cohen–Macaulay formal fibers” via the argument already present in [Lyu-dual-complex-lift, §8].
Finding the assignment for the lci property is the most difficult. This is because the intrinsic invariant that determines a Noetherian ring $A$ is lci or not, namely the cotangent complex $L_{A/\mathbf{Z}}$, does not have finite cohomology modules. When $A$ is complete local, we can find a regular local ring $R$ and a surjective map $R\to A$, and $L_{A/R}$ does have finite cohomology modules; however, this is still fragile with respect to ring maps. In any case, we need a way to use an ideal to detect the non-flatness of certain non-finite modules, more explicitly the modules $C_{n}$ appearing in [Briggs-Iyenger-Cotangent-Complex]. We investigate their structures in §11, and successfully define their Fitting invariant (for $n\geq\dim A+2$), which does the trick. These observations, namely the structure of $C_{n}$ (Lemma 11.3) and Fitting invariant for certain non-finite modules (Definition 11.6), may be of their own interest.
@@
Acknowledgment. The author thanks Pham Hung Quy for suggesting the author to consider the lifting problem for CM-quotients. The author thanks Kevin Tucker and Wenliang Zhang for helpful discussions. The author was supported by an AMS-Simons Travel Grant.
@@ 3. Subalgebras of normalization
See [Nagata-local, §33] for relevant materials.
@@
Let $R$ be a Noetherian integral domain, $S$ a sub-$R$-algebra of $R^{\nu}$. Then for every $\mathfrak{p}\in\operatorname{Spec}(R)$, there are only finitely many $\mathfrak{q}\in\operatorname{Spec}(S)$ above $\mathfrak{p}$, and $\kappa(\mathfrak{q})$ is finite over $\kappa(\mathfrak{p})$ for all $\mathfrak{q}$.
@@
If $S=R^{\nu}$, then this is part of [Nagata-local, Theorem 33.10]. The general case follows from the fact $\operatorname{Spec}(R^{\nu})\to\operatorname{Spec}(S)$ is surjective. ∎
Lemma 3.2 (cf. [Nagata-local, (33.11)]).
Let $R$ be a Noetherian integral domain, $S$ a sub-$R$-algebra of $R^{\nu}$. Let $a\in S^{\circ}$, and let $\mathfrak{q}$ be a minimal prime divisor of $aS$.
@@ 1. $(i)$
    There exists a finite $R$-subalgebra $R^{\prime}$ of $S$ such that $\operatorname{ht}(\mathfrak{q}\cap R^{\prime})=1$.
@@ 2. $(ii)$
Let $R$ be a Noetherian integral domain, $S$ a sub-$R$-algebra of $R^{\nu}$. Then for every $a\in S^{\circ}$, the set of minimal prime divisors $\mathfrak{q}$ of $aS$ is finite, and for every $\mathfrak{q}$, $S_{\mathfrak{q}}$ is a $1$-dimensional Noetherian ring.
@@
We may assume $a\in R$. Finiteness follows from Lemma 3.2$(ii)$ and Theorem 3.1. For each $\mathfrak{q}$, Lemma 3.2$(i)$ gives a map $R^{\prime}_{\mathfrak{q}\cap R^{\prime}}\to S_{\mathfrak{q}}$, so $S_{\mathfrak{q}}$ is a $1$-dimensional Noetherian ring by the theorem of Krull-Akizuki [Nagata-local, Theorem 33.2]. ∎
@@
Let $R$ be a Noetherian integral domain, $S$ a sub-$R$-algebra of $R^{\nu}$. We say $S$ is $(S_{2})$ if for all $a\in S^{\circ}$, $aS$ is a finite intersection of primary ideals of height $1$. This is the same as Serre’s condition $(S_{2})$ if $S$ is Noetherian.
@@
Let $R$ be a Noetherian integral domain, $S$ a sub-$R$-algebra of $R^{\nu}$. Then the following are equivalent.
@@ 3. $(iii)$
That $(i)$ implies $(ii)$ is clear as minimal prime divisors of and the set of zero divisors modulo $aS$ can be read off of a primary decomposition, cf. [AMcommalg, Proposition 4.7].
@@
Let $R$ be a Noetherian integral domain, $S$ a sub-$R$-algebra of $R^{\nu}$. Then the following are true.
@@ 2. $(ii)$
Let $R$ be a Noetherian integral domain, $S$ a sub-$R$-algebra of $R^{\nu}$. Assume that $S$ is the filtered union of $R$-subalgebras $\{S_{\alpha}\}_{\alpha}$, and that each $S_{\alpha}$ is $(S_{2})$. Then $S$ is $(S_{2})$.
@@ 4. Naive and canonical $(S_{2})$-closures
For a Noetherian integral domain $R$ and a sub-$R$-algebra $S$ of $R^{\nu}$, we know $S$ is $(S_{2})$ if and only if $S=S^{n\sigma}$ (Lemma 3.5). We will see $S^{\sigma}$ is $(S_{2})$ (Theorem 6.8), so $S$ is $(S_{2})$ if and only if $S=S^{\sigma}$.
@@
Let $R$ be a Noetherian integral domain, $S$ a sub-$R$-algebra of $R^{\nu}$.
@@
Let $S$ be an integral domain, Let $S^{\prime}$ be a sub-$S$-algebra of $S^{n\sigma}$ (resp. $S^{\sigma}$). Then $S^{\prime n\sigma}\subseteq S^{n\sigma}$ (resp. $S^{\prime\sigma}\subseteq S^{\sigma}$).
@@
By a theorem of Lech [Lech-completion-domain], a complete Noetherian local ring containing a field is the completion of a Noetherian local domain if and only if its depth is at least $1$. Therefore, there exist a Noetherian local domain $R$ of dimension $2$ whose completion is $k[[x,y,z]]/(x,y)\cap(z)$, where $k$ is a field.
@@
If the field $k$ has characteristic zero, then we can even make $R$ quasi-excellent; in fact, a Noetherian complete local ring containing a field of characteristic zero is the completion of a quasi-excellent local domain if and only if it is reduced. See [Loepp-03-complete-excellent-domains, Proof of Theorem 9].
@@
Assume $R$ is semilocal. For $P\in\operatorname{OS}_{2}^{\wedge}({R})$, the punctured spectrum of the ring $(R^{\wedge})_{P}$ is disconnected, as it has an isolated point given by a minimal prime $P_{0}\subsetneq P$ with $\operatorname{ht}(P/P_{0})=1$. By [stacks, Tag 0BLR] we have $20pt(R^{\wedge})_{P}<$ 2, so $20ptR_{P\cap R}<$ 2. Therefore, if $R$ is $(S_{2})$, then $\operatorname{OS}_{2}^{\wedge}({R})=\emptyset$.
@@ 3. $(iii)$
    If a sub-$R$-algebra $S$ of $R^{\sigma}$ is $(S_{2})$, then $S=R^{\sigma}$.
@@ 4. $(iv)$
Let $0=Q_{1}\cap\ldots\cap Q_{m}\cap Q_{m+1}\cap\ldots\cap Q_{n}$ be a shortest primary decomposition of 00 in $R^{\wedge}$, ordered in a way that $P_{j}=\sqrt{Q_{j}}$ is a minimal prime of $R^{\wedge}$ for $1\leq j\leq m$ and not for $m+1\leq j\leq n$. Let $S_{j}=R^{\wedge}/Q_{j}$ for $1\leq j\leq m$, $S=\prod_{j=1}^{m}S_{j}$. Let $L_{j}=(S_{j})_{P_{j}}$, so $L=\prod_{j=1}^{m}L_{j}$ is the total fraction ring of $S$.
@@
[…]
@@
Observe that for every $\mathfrak{p}\in\operatorname{Spec}_{1}(R)$, we have $R_{\mathfrak{p}}=(R^{n\sigma})_{\mathfrak{p}}$. As $\operatorname{OS}_{2}^{\wedge}({R})=\emptyset$, we see that for every $P\in\operatorname{Spec}(R^{\wedge})$ so that $\operatorname{ht}(P/P_{j})=1$ for some $1\leq j\leq m$, we have $(R^{\wedge})_{P}=(R^{n\sigma}\otimes_{R}R^{\wedge})_{P}$. Therefore the sub-$S_{j}$-algebra $T_{j}$ of $L_{j}$ satisfies $(S_{j})_{P}=(T_{j})_{P}$ for all $P\in\operatorname{Spec}_{1}(S_{j})$.
@@
Let $A$ be a Noetherian complete local ring that is $(S_{1})$ and has irreducible spectrum. Let $F$ be the total fraction ring of $A$ and let $M\subseteq N$ be two sub-$A$-modules of $F$. If $M_{P}=N_{P}$ for all $P\in\operatorname{Spec}_{1}(A)$, and $M$ is finite, then $N$ is finite.
@@
Let $U$ be the locus where $M$ is $(S_{2})$. Then $U$ is open [EGA4_2, Proposition 6.11.6], and contains $\operatorname{Spec}_{1}(A)$ as $A$ is $(S_{1})$. Let $j:U\to\operatorname{Spec}(A)$ be the canonical open immersion. It is clear that $j_{*}j^{*}M=\bigcap_{P\in\operatorname{Spec}_{1}(A)}M_{P}$.
Let $P_{0}$ be the minimal prime of $A$, and let $\overline{A}=A/P_{0}$. Note that $\operatorname{Ass}_{A}(M)=\{P_{0}\}$. For $P\in\operatorname{Spec}(A)\setminus U$, the completion $\overline{A}^{\wedge}_{P}$ is $(S_{1})$ as $\overline{A}$ is $(S_{1})$ with $(S_{1})$ formal fibers. Moreover, by Ratliff [stacks, Tag 0AW6], $\overline{A}^{\wedge}_{P}$ is equidimensional, as $\overline{A}_{P}$ is equidimensional and universally catenary. Therefore for all $\mathfrak{p}\in\operatorname{Ass}(\overline{A}^{\wedge}_{P})$ we have $\dim\overline{A}^{\wedge}_{P}/\mathfrak{p}=\operatorname{ht}(P)>1$. By [stacks, Tag 0BK3], $j_{*}j^{*}M$ is finite; and so is its submodule $N$. ∎
@@
Let $S$ be an arbitrary sub-$R$-algebra of $R^{n\sigma}=R^{\sigma}$. Then $S^{n\sigma}=R^{n\sigma}$ by Lemmas 4.3 and 4.10 below. Therefore $S$ is $(S_{2})$ if and only if $S=R^{n\sigma}$. This gives $(ii)$$(iii)$. ∎
@@
Let $R\subseteq R^{\prime}$ be a finite inclusion of Noetherian integral domains, $S^{\prime}$ a sub-$R^{\prime}$-algebra of $R^{\prime\nu}$. Then for every $\mathfrak{q}^{\prime}\in\operatorname{Spec}_{1}(S^{\prime})$, we have either $\mathfrak{q}^{\prime}\cap R\in\operatorname{Spec}_{1}(R)$, or $\mathfrak{q}^{\prime}\cap R=P\cap R$ for some $P\in\operatorname{OS}_{2}^{\wedge}({R})$; in particular, if $\operatorname{OS}_{2}^{\wedge}({R})=\emptyset$, then $\mathfrak{q}^{\prime}\cap R\in\operatorname{Spec}_{1}(R)$, and $R^{n\sigma}\subseteq S^{\prime n\sigma}$.
@@
Let $R$ be a Noetherian integral domain, $S$ a sub-$R$-algebra of $R^{\nu}$. Assume that $\operatorname{OS}_{2}^{\wedge}({R})=\emptyset$. Then the following are true.

### v2

@@ 1. Introduction
When Question 1.1 admits an affirmative answer, we say $\mathbf{R}$ has the lifting property. There have been numerous studies on the lifting problem and its variants, for many important properties $\mathbf{R}$. For example, lifting holds for $\mathbf{R}$=“Nagata” [Mar75] and $\mathbf{R}$=“quasi-excellent” [KS21], but not for $\mathbf{R}$=“excellent” or $\mathbf{R}$=“universally catenary” [Gre82]. We refer the reader to [KS21, Appendix] for more information.
In previous work (see [Lyu, §8]), the author showed that $\mathbf{R}$=“is a quotient of a Gorenstein ring” satisfies the lifting property. However, whether or not $\mathbf{R}$=“is a quotient of a Cohen–Macaulay ring” satisfies the lifting property seems to be difficult. We provide two perspectives on this question.
@@
We call a ring $R$ semi-Nagata if $R$ is Noetherian and for every finite $R$-algebra $B$ that is an integral domain, $B$ admits a finite $(S_{2})$-ification, in the sense that there is a finite inclusion of integral domains $B\to C$ so that $C$ is $(S_{2})$ and $B_{\mathfrak{p}}=C_{\mathfrak{p}}$ for all $\mathfrak{p}\in\operatorname{Spec}_{1}(B)$. We have the following main result, which the author believes to be new (Definition 7.1 and Theorems 7.8 and 7.14)
@@ 4. $(iv)$
See [EGA ̵IV2, §7.6 and §7.7] for (i)(i)(ii)(ii)$(iii)$, whereas $(iv)$ is trivial. In particular, from either $(iii)$ or $(iv)$ of both theorems, we have the following result, which the author also believes to be new.
@@
There are DVRs that are not Nagata, [Nag62, Appendix A1, Example 3]. On the other hand we have (Remark 7.2 and Corollary 7.15)
@@
To show $R^{\sigma}$ is $(S_{2})$, and to show our main Theorem 1.2, we show that for a semilocal $R$, there exists a finite subalgebra of $R^{\sigma}$ that do not have FONSIs (Theorem 6.4). After all, FONSIs are pretty rare (Remarks 4.6 and 4.7 and Lemma 6.3). The idea is inspired by a classical argument of Ratliff [Mat89, §31, Lemma 4]. We use local cohomology to make a conceptual argument.
@@
To conclude the first part, we show that lifting of the semi-Nagata property holds for universally catenary rings.
Theorem 1.6 (=Theorem 8.1).
Let $R$ be a Noetherian ring, $I$ an ideal of $R$. Assume that
  1. (1)
    $R$ is $I$-adically complete.
  2. (2)
    $R/I$ is semi-Nagata.
  3. (3)
    $R$ is universally catenary.
Then $R$ is semi-Nagata.
The author was not able to show lifting in full generality. However, when restricted to semilocal rings, much more advances in the lifting problem are made in the second part (§§913) of this article. We repeat the problem in this setting.
Question 1.7 (local lifting problem).
@@
Question 1.7 admits an affirmative answer when $\mathbf{R}$=“has $(S_{k})$ formal fibers,” where $k\geq 0$ is arbitrary, “has Cohen–Macaulay formal fibers,” “has Gorenstein formal fibers,” “has lci formal fibers,” and “is a quotient of a Cohen–Macaulay ring.”
Note that when $k=1$, this is exactly the semi-Nagata property (Theorem 1.2(i)(i)).
We show the formal fiber properties in Theorem 1.8 following the argument of Nishimura [Nis81]. The key new input is that we find ideals that define the non-$\mathbf{P}$-locus (where $\mathbf{P}$$(S_{k})$,”“Cohen–Macaulay,” “Gorenstein,” or “lci”) of nice rings strictly functorially with respect to nice homomorphisms. This task was effortless for the properties considered in [Nis81]. We explain this in §9.
In our case, such an assignment of ideals can be found with effort. We do a basic reduction in §10, saying we just need to assign $\mathfrak{m}$-primary ideals to complete local rings $(A,\mathfrak{m})$ that are $\mathbf{P}$ exactly on the punctured spectrum, strictly functorial with respect to flat maps with $\mathbf{P}$-fibers and 0-dimensional special fiber. After the reduction we find a desired assignment for all properties but lci in §11. The case $\mathbf{R}$=“is a quotient of a Cohen–Macaulay ring” of Theorem 1.8 follows from the case $\mathbf{R}$=“has Cohen–Macaulay formal fibers” via the argument already present in [Lyu, §8].
Finding the assignment for the lci property is the most difficult. This is because the intrinsic invariant that determines a Noetherian ring $A$ is lci or not, namely the cotangent complex $L_{A/\mathbf{Z}}$, does not have finite cohomology modules. When $A$ is complete local, we can find a regular local ring $R$ and a surjective map $R\to A$, and $L_{A/R}$ does have finite cohomology modules; however, this is still fragile with respect to ring maps. In any case, we need a way to use an ideal to detect the non-flatness of certain non-finite modules, more explicitly the modules $C_{n}$ appearing in [BI23]. We investigate their structures in §12, and successfully define their Fitting invariant (for $n\geq\dim A+2$), which does the trick. These observations, namely the structure of $C_{n}$ (Lemma 12.3) and Fitting invariant for certain non-finite modules (Definition 12.6), may be of their own interest.
@@
Acknowledgment. The author thanks Pham Hung Quy for suggesting the author to consider the lifting problem for CM-quotients. The author thanks Linquan Ma, Kevin Tucker, and Wenliang Zhang for helpful discussions. The author was supported by an AMS-Simons Travel Grant.
@@ 3. Subalgebras of normalization
See [Nag62, §33] for relevant materials.
@@
Let $R$ be a Noetherian integral domain, $S$ a subalgebra of $R^{\nu}$. Then for every $\mathfrak{p}\in\operatorname{Spec}(R)$, there are only finitely many $\mathfrak{q}\in\operatorname{Spec}(S)$ above $\mathfrak{p}$, and $\kappa(\mathfrak{q})$ is finite over $\kappa(\mathfrak{p})$ for all $\mathfrak{q}$.
@@
If $S=R^{\nu}$, then this is part of [Nag62, Theorem 33.10]. The general case follows from the fact $\operatorname{Spec}(R^{\nu})\to\operatorname{Spec}(S)$ is surjective. ∎
Lemma 3.2 (cf. [Nag62, (33.11)]).
Let $R$ be a Noetherian integral domain, $S$ a subalgebra of $R^{\nu}$. Let $a\in S^{\circ}$, and let $\mathfrak{q}$ be a minimal prime divisor of $aS$.
@@ 1. $(i)$
    There exists a finite subalgebra $R^{\prime}$ of $S$ such that $\operatorname{ht}(\mathfrak{q}\cap R^{\prime})=1$.
@@ 2. $(ii)$
Let $R$ be a Noetherian integral domain, $S$ a subalgebra of $R^{\nu}$. Then for every $a\in S^{\circ}$, the set of minimal prime divisors $\mathfrak{q}$ of $aS$ is finite, and for every $\mathfrak{q}$, $S_{\mathfrak{q}}$ is a $1$-dimensional Noetherian ring.
@@
We may assume $a\in R$. Finiteness follows from Lemma 3.2(ii)(ii) and Theorem 3.1. For each $\mathfrak{q}$, Lemma 3.2(i)(i) gives a map $R^{\prime}_{\mathfrak{q}\cap R^{\prime}}\to S_{\mathfrak{q}}$, so $S_{\mathfrak{q}}$ is a $1$-dimensional Noetherian ring by the theorem of Krull–Akizuki [Nag62, Theorem 33.2]. ∎
@@
Let $R$ be a Noetherian integral domain, $S$ a subalgebra of $R^{\nu}$. We say $S$ is $(S_{2})$ if for all $a\in S^{\circ}$, $aS$ is a finite intersection of primary ideals of height $1$. This is the same as Serre’s condition $(S_{2})$ if $S$ is Noetherian.
@@
Let $R$ be a Noetherian integral domain, $S$ a subalgebra of $R^{\nu}$. Then the following are equivalent.
@@ 3. $(iii)$
That $(i)$ implies $(ii)$ is clear as minimal prime divisors of and the set of zero divisors modulo $aS$ can be read off of a primary decomposition, cf. [AM69, Proposition 4.7].
@@
Let $R$ be a Noetherian integral domain, $S$ a subalgebra of $R^{\nu}$. Then the following are true.
@@ 2. $(ii)$
Let $R$ be a Noetherian integral domain, $S$ a subalgebra of $R^{\nu}$. Assume that $S$ is the filtered union of $R$-subalgebras $\{S_{\alpha}\}_{\alpha}$, and that each $S_{\alpha}$ is $(S_{2})$. Then $S$ is $(S_{2})$.
@@ 4. Naive and canonical $(S_{2})$-closures
For a Noetherian integral domain $R$ and a subalgebra $S$ of $R^{\nu}$, we know $S$ is $(S_{2})$ if and only if $S=S^{n\sigma}$ (Lemma 3.5). We will see $S^{\sigma}$ is $(S_{2})$ (Theorem 6.8), so $S$ is $(S_{2})$ if and only if $S=S^{\sigma}$.
@@
Let $R$ be a Noetherian integral domain, $S$ a subalgebra of $R^{\nu}$.
@@
Let $S$ be an integral domain, Let $S^{\prime}$ be a subalgebra of $S^{n\sigma}$ (resp. $S^{\sigma}$). Then $S^{\prime n\sigma}\subseteq S^{n\sigma}$ (resp. $S^{\prime\sigma}\subseteq S^{\sigma}$).
@@
By a theorem of Lech [Lec86], a complete Noetherian local ring containing a field is the completion of a Noetherian local domain if and only if its depth is at least $1$. Therefore, there exist a Noetherian local domain $R$ of dimension $2$ whose completion is $k[[x,y,z]]/(x,y)\cap(z)$, where $k$ is a field.
@@
If the field $k$ has characteristic zero, then we can even make $R$ quasi-excellent; in fact, a Noetherian complete local ring containing a field of characteristic zero is the completion of a quasi-excellent local domain if and only if it is reduced. See [Loe03, Proof of Theorem 9].
@@
Assume $R$ is semilocal. For $P\in\operatorname{OS}_{2}^{\wedge}({R})$, the punctured spectrum of the ring $(R^{\wedge})_{\mathfrak{P}}$ is disconnected, as it has an isolated point given by a minimal prime $P_{0}\subsetneq P$ with $\operatorname{ht}(P/P_{0})=1$. By [Stacks, Tag 0BLR] we have $0pt(R^{\wedge})_{\mathfrak{P}}<2$, so $0ptR_{P\cap R}<2$. Therefore, if $R$ is $(S_{2})$, then $\operatorname{OS}_{2}^{\wedge}({R})=\emptyset$.
@@ 3. $(iii)$
    If a subalgebra $S$ of $R^{\sigma}$ is $(S_{2})$, then $S=R^{\sigma}$.
@@ 4. $(iv)$
Let $0=Q_{1}\cap\ldots\cap Q_{m}\cap Q_{m+1}\cap\ldots\cap Q_{n}$ be a shortest primary decomposition of 0 in $R^{\wedge}$, ordered in a way that $P_{j}=\sqrt{Q_{j}}$ is a minimal prime of $R^{\wedge}$ for $1\leq j\leq m$ and not for $m+1\leq j\leq n$. Let $S_{j}=R^{\wedge}/Q_{j}$ for $1\leq j\leq m$, $S=\prod_{j=1}^{m}S_{j}$. Let $L_{j}=(S_{j})_{P_{j}}$, so $L=\prod_{j=1}^{m}L_{j}$ is the total fraction ring of $S$.
@@
$R^{\wedge}$$R^{n\sigma}\otimes_{R}R^{\wedge}$$K\otimes_{R}R^{\wedge}$}S{S}T{T}L{L}
@@
Observe that for every $\mathfrak{p}\in\operatorname{Spec}_{1}(R)$, we have $R_{\mathfrak{p}}=(R^{n\sigma})_{\mathfrak{p}}$. As $\operatorname{OS}_{2}^{\wedge}({R})=\emptyset$, we see that for every $P\in\operatorname{Spec}(R^{\wedge})$ so that $\operatorname{ht}(P/P_{j})=1$ for some $1\leq j\leq m$, we have $(R^{\wedge})_{\mathfrak{P}}=(R^{n\sigma}\otimes_{R}R^{\wedge})_{\mathfrak{P}}$. Therefore the sub-$S_{j}$-algebra $T_{j}$ of $L_{j}$ satisfies $(S_{j})_{P}=(T_{j})_{\mathfrak{P}}$ for all $P\in\operatorname{Spec}_{1}(S_{j})$.
@@
Let $A$ be a Noetherian complete local ring that is $(S_{1})$ and has irreducible spectrum. Let $F$ be the total fraction ring of $A$ and let $M\subseteq N$ be two sub-$A$-modules of $F$. If $M_{P}=N_{\mathfrak{P}}$ for all $P\in\operatorname{Spec}_{1}(A)$, and $M$ is finite, then $N$ is finite.
@@
Let $U$ be the locus where $M$ is $(S_{2})$. Then $U$ is open [EGA ̵IV2, Proposition 6.11.6], and contains $\operatorname{Spec}_{1}(A)$ as $A$ is $(S_{1})$. Let $j:U\to\operatorname{Spec}(A)$ be the canonical open immersion. It is clear that $j_{*}j^{*}M=\bigcap_{P\in\operatorname{Spec}_{1}(A)}M_{\mathfrak{P}}$.
Let $P_{0}$ be the minimal prime of $A$, and let $\overline{A}=A/P_{0}$. Note that $\operatorname{Ass}_{A}(M)=\{P_{0}\}$. For $P\in\operatorname{Spec}(A)\setminus U$, the completion $\overline{A}^{\wedge}_{\mathfrak{P}}$ is $(S_{1})$ as $\overline{A}$ is $(S_{1})$ with $(S_{1})$ formal fibers. Moreover, by Ratliff [Stacks, Tag 0AW6], $\overline{A}^{\wedge}_{\mathfrak{P}}$ is equidimensional, as $\overline{A}_{\mathfrak{P}}$ is equidimensional and universally catenary. Therefore for all $\mathfrak{p}\in\operatorname{Ass}(\overline{A}^{\wedge}_{\mathfrak{P}})$ we have $\dim\overline{A}^{\wedge}_{\mathfrak{P}}/\mathfrak{p}=\operatorname{ht}(P)>1$. By [Stacks, Tag 0BK3], $j_{*}j^{*}M$ is finite; and so is its submodule $N$. ∎
@@
Let $S$ be an arbitrary subalgebra of $R^{n\sigma}=R^{\sigma}$. Then $S^{n\sigma}=R^{n\sigma}$ by Lemmas 4.3 and 4.10 below. Therefore $S$ is $(S_{2})$ if and only if $S=R^{n\sigma}$. This gives (ii)(ii)$(iii)$. ∎
@@
Let $R\subseteq R^{\prime}$ be a finite inclusion of Noetherian integral domains, $S^{\prime}$ a subalgebra of $R^{\prime\nu}$. Then for every $\mathfrak{q}^{\prime}\in\operatorname{Spec}_{1}(S^{\prime})$, we have either $\mathfrak{q}^{\prime}\cap R\in\operatorname{Spec}_{1}(R)$, or $\mathfrak{q}^{\prime}\cap R=P\cap R$ for some $P\in\operatorname{OS}_{2}^{\wedge}({R})$; in particular, if $\operatorname{OS}_{2}^{\wedge}({R})=\emptyset$, then $\mathfrak{q}^{\prime}\cap R\in\operatorname{Spec}_{1}(R)$, and $R^{n\sigma}\subseteq S^{\prime n\sigma}$.
@@
Let $R$ be a Noetherian integral domain, $S$ a subalgebra of $R^{\nu}$. Assume that $\operatorname{OS}_{2}^{\wedge}({R})=\emptyset$. Then the following are true.